OpenStudy (anonymous):

The measures of two base angles of an isosceles trapezoid are 7x – 12 and 5x + 6. What is the measure of x? Select one: a. 3 b. 5 c. 7 d. 9

OpenStudy (some.random.cool.kid):

7x - 12 = 5x + 6 -5x+12 -5x +12 2x = 18 /2 /2 x = 9

OpenStudy (some.random.cool.kid):

@LTM

OpenStudy (anonymous):

Thanks (: i dont know geometry whatso ever